OpenStudy (anonymous):

A jar of mayonnaise is shaped like a cylinder. The radius of the jar is 1.5 in. and its height is 7 in. The mayonnaise is on sale for $0.08 per in3. What is the cost of the jar of mayonnaise, rounded to the nearest cent? Use 3.14 to approximate pi. $

OpenStudy (anonymous):

Find the volume of the cylinder. \(V = \pi r^2 h\) \(V = (3.14)(1.5^2)(7)\)

OpenStudy (anonymous):

Simplify.

OpenStudy (anonymous):

49.455

OpenStudy (anonymous):

Correct, now multiply by 0.08

OpenStudy (anonymous):

3.95?

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

That is correct, that's our answer.

OpenStudy (anonymous):

Actually, if you want to round it should be 3.96

OpenStudy (anonymous):

oh it wouldn't be 4.00?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

It says to the nearest cent, not dollar.